# X-Vine Models for Multivariate Extremes

### v1

@@ B.5 Proofs for Section 6
Next, let $d\geqslant 4$ and assume the statement is true for regular vine sequence on $d-1$ elements. Let $e=(a_{e},b_{e};D_{e})=\left\{{f,g}\right\}$ be the single edge in $E_{d-1}$, with $f,g\in E_{d-2}$ labeled in such a way that $a_{e}\in\mathcal{C}_{f}$ and $b_{e}\in\mathcal{C}_{g}$. The regular vine sequence $\mathcal{V}$ induces regular vine sequences $\mathcal{V}_{f}$ and $\mathcal{V}_{g}$ on the node sets $A_{f}=[d]\setminus\left\{{b_{e}}\right\}$ and $A_{g}=[d]\setminus\left\{{a_{e}}\right\}$, both of which have $d-1$ elements; the edges $\bar{e}$ of the two regular vine sequences $\mathcal{V}_{f}$ and $\mathcal{V}_{g}$ are those of $\mathcal{V}$ that satisfy $A_{\bar{e}}\subseteq A_{f}$ or $A_{\bar{e}}\subseteq A_{g}$, respectively. Let $j\in[d]$. If $j\neq b_{e}$, then we can apply the induction hypothesis to $\mathcal{V}_{f}$ to find a bijection $\sigma_{f,j}:\left\{{1,\ldots,d-1}\right\}\to A_{f}=[d]\setminus\left\{{b_{e}}\right\}$ such that (i) $\sigma_{f,j}(1)=j$ and (ii) $\sigma_{f,j}(k)$ belongs to the conditioned set of some edge $e_{k-1}$ in the $(k-1)$th tree of $\mathcal{V}_{f}$ as well as $\left\{{\sigma_{f,j}(i):i=1,\ldots,k}\right\}=A_{e_{k-1}}$, for all $k\in\left\{{2,\ldots,d-1}\right\}$. Extend $\sigma_{f,j}$ to a map $\sigma_{j}$ on $[d]$ by $\sigma_{j}(d)=b_{e}$. Then $\sigma_{j}$ meets all the requirements. If $j=b_{e}$, then $j\neq a_{e}$, and we can apply the same argument to $\mathcal{V}_{g}$. ∎
@@ B.6 Proofs for Section 7
(i) By Definition 6.1, the density of $\bm{Z}$ is the function $\bm{z}\mapsto r(\bm{z})\operatorname{\mathds{1}}_{\mathbb{L}}(\bm{z})/R(\mathbb{L})$ with $\mathbb{L}=\left\{{\bm{x}:\min(\bm{z})<1}\right\}$. Let $K\subset[d]$ be non-empty and let $\bm{z}_{K}\in(0,\infty)^{K}$ be such that $\min\bm{z}<1$. The marginal density of $\bm{Z}_{K}=(Z_{k})_{k\in K}$ evaluated in $\bm{z}_{K}$ is
@@
Since $\min\bm{z}_{K}<1$, we also have $\min\bm{z}<1$ for $\bm{z}_{[d]\setminus K}\in(0,\infty)^{[d]\setminus K}$. As a consequence, the indicator function in the integral is always equal to one and, by Eq. (2.7), the density of $\bm{Z}_{K}$ evaluated in $\bm{z}_{K}$ simplifies to $r_{K}(\bm{z}_{K})/r(\mathbb{L})$. Applying this identity to $K=J$ and $K=I\cup J$, we find, for $\bm{z}_{I}\in(0,\infty)^{I}$ and for $\bm{z}_{J}\in(0,\infty)^{J}$ such that $\min\bm{z}_{J}<1$ and $r_{J}(\bm{z}_{J})>0$, the conditional density of $\bm{Z}_{I}$ given $\bm{Z}_{J}=\bm{z}_{J}$ evaluated in $\bm{z}_{I}$:
@@
Appendix C Converse of Sklar’s theorem for tail copula densities
As a converse to Sklar’s theorem for tail copula densities (Proposition 3.1), we can combine several tail copula densities along a scale-invariant family of copula densities to form a tail copula density in higher dimensions. With slight abuse of notation, write $i\cup J:=\{i\}\cup J$ for $i\in I$. A family of $k$-variate copula densities $\left\{{c(\,\cdot\,;\theta):\theta\in\Theta}\right\}$ indexed by a Borel set $\Theta\subseteq\mathbb{R}^{m}$ is jointly measurable if the map $(\bm{u},\theta)\mapsto c(\bm{u};\theta)$ is Borel measurable on $(0,1)^{k}\times\Theta$. Joint measurability is trivial if $c(\,\cdot\,;\theta)$ does not depend on $\theta$.
Proposition C.1 (Sklar’s theorem: converse part).
$r(\bm{x}):=r_{J}(\bm{x}_{J})\cdot\prod_{i\in I}r_{i|J}(x_{i}|\bm{x}_{J})\cdot c_{I;J}\left({R_{i|J}(x_{i}|\bm{x}_{J}),i\in I;\bm{x}_{J}}\right),$ (C.1)
for $\bm{x}\in(0,\infty)^{d}$ such that $r_{J}(\bm{x}_{J})>0$ and zero otherwise is a $d$-variate tail copula density with margins $r_{i\cup J}$ for $i\in I$.
Proof of Proposition C.1.
First, we show that $r$ is a tail copula density. To this end, we need to show two properties: the normalization constraint (2.4) and the homogeneity relation (2.5). To show the homogeneity relation (2.5), let $t\in(0,\infty)$, and let $\bm{x}\in(0,\infty)^{d}$ be such that $r_{J}(\bm{x}_{J})>0$. Then
$(1-|J|)(1-|I|)-|I|\cdot|J|=1-|I|-|J|=1-d.$
To show the normalization constraint (2.4), let $k\in\left\{{1,\ldots,d}\right\}$ and $x_{k}\in(0,\infty)$. We need to show that $\int_{(0,\infty)^{d-1}}r(\bm{x})\,\mathrm{d}\bm{x}_{\setminus k}=1$, where the integral is over all coordinates except for the $k$th one. There are two cases: $k\in I$ or $k\in J$. First, if $k\in I$, the integral of interest is
as $r_{k\cup J}$ is a tail copula density. Second, if $k\in J$, the integral of interest is
since $r_{J}$ is a tail copula density. The proof that $r$ has the required margins $r_{i\cup J}$ for $i\in I$ is similar to the case $k\in I$ in the previous part of the proof. Finally, statement (3.1) easily follows by applying Sklar’s theorem to $r_{I|J}(\,\cdot\,|\bm{x}_{J})$. ∎
In the special case that $J$ is a singleton, $J=\left\{{j}\right\}$, we have $I=[d]\setminus j$, the $(d-1)$ tail copula densities $(r_{i,j})_{i\in[d]\setminus j}$ are bivariate, and the common-margin assumption is automatic by (2.2); in particular, $r_{i|j}(x_{i}|x_{j})=r_{i,j}(x_{i},x_{j})$. Further, the scale-invariance property (3.2) enforces the simplifying assumption, so that the construction (C.1) becomes
Appendix D Exponent measure densities as X-vines
@@
$c_{I;J}(\bm{u}_{I};1/\bm{y}_{J})=\frac{\lambda(\bm{y}_{I}|\bm{y}_{J})}{\prod_{i\in I}\lambda_{i|J}(y_{i}|\bm{y}_{J})}\quad\text{with}\quad y_{i}=\overline{\Lambda}_{i|J}^{-1}(u_{i}|\bm{y}_{J}).$ (D.1)
@@
Appendix E X-vines as limits of regular vine copula densities
The aim of this section is to link up regular vine copula densities as introduced in Bedford and Cooke (2002) with X-vine tail copula densities, as claimed in Remark 5.8. To this end, we start from a $d$-variate regular vine copula density $c$ and we show that it is attracted to a $d$-variate X-vine tail copula density $r$ along exactly the same regular vine sequence as $c$. Moreover, $c$ and $r$ share the same pair copula densities on all edges except for those in the first tree. For edges in the first tree, the pair copula densities of $c$ are to be replaced by the bivariate tail copula densities to which they are attracted.
$c(\bm{u})=\prod_{e\in E_{1}}c_{a_{e},b_{e}}(u_{a_{e}},u_{b_{e}})\cdot\prod_{j=2}^{d-1}\prod_{e\in E_{j}}c_{a_{e},b_{e};D_{e}}\left({C_{a_{e}|D_{e}}\left({u_{a_{e}}|\bm{u}_{D_{e}}}\right),\,C_{b_{e}|D_{e}}\left({u_{b_{e}}|\bm{u}_{D_{e}}}\right)}\right),$ (E.1)
$C_{a_{e},b_{e}|D_{e}}\left({u_{a_{e}},u_{b_{e}}|\bm{u}_{D_{e}}}\right)=C_{a_{e},b_{e};D_{e}}\left({C_{a_{e}|D_{e}}\left({u_{a_{e}}|\bm{u}_{D_{e}}}\right),\,C_{b_{e}|D_{e}}\left({u_{b_{e}}|\bm{u}_{D_{e}}}\right)}\right).$ (E.2)
Assumption E.1.
Let $c$ be a $d$-variate regular vine copula density as in (E.1), with $d\geqslant 3$.
  1. (i)
    For every edge $e\in E_{1}$, there exists a bivariate tail copula density $r_{a_{e},b_{e}}$ such that
    $\lim_{t\searrow 0}t\cdot c_{a_{e},b_{e}}(tx,ty)=r_{a_{e},b_{e}}(x,y),\qquad(x,y)\in(0,\infty)^{2}.$
  2. (ii)
Proposition E.2.
Under Assumption E.1, we have
$\lim_{t\searrow 0}t^{d-1}\cdot c(t\bm{x})=r(\bm{x}),\qquad\bm{x}\in(0,\infty)^{d},$
where the X-vine tail copula density $r$ is generated by the triple $(\mathcal{V},\mathcal{R},\mathcal{C})$ as in Theorem 5.7 with
  • $\mathcal{V}$ the same regular vine sequence as in (E.1),
  • =$\mathcal{R}=\left\{{r_{a_{e},b_{e}}:e\in E_{1}}\right\}$ for $r_{a_{e},b_{e}}$ in Assumption E.1(i), and
  • $\mathcal{C}$ the same bivariate copula densities for edges $e\in E_{j}$ with $j\geqslant 2$ as in (E.1).
Proof of Proposition E.2.
Let $\bm{x}\in(0,\infty)^{d}$. As there $d-1$ edges in the first tree, we have
The first product on the right-hand side converges to $\prod_{e\in E_{1}}r_{a_{e},b_{e}}(x_{a_{e}},x_{b_{e}})$ as $t\to 0$ by Assumption E.1(i). It remains to show that for every $j\in\left\{{2,\ldots,d-1}\right\}$, every $e\in E_{j}$ and every $i\in\mathcal{C}_{e}=\left\{{a_{e},b_{e}}\right\}$, we have
$\lim_{t\searrow 0}C_{i|D_{e}}\left({tx_{i}|t\bm{x}_{D_{e}}}\right)=R_{i|D_{e}}\left({x_{i}|\bm{x}_{D_{e}}}\right).$ (E.3)
By properties of regular vine sequences as explained in Section 2.2, there exists an edge $f=(a_{f},b_{f};D_{f})\in E_{j-1}$ such that $f\in e$ and $i\in\mathcal{C}_{f}=\left\{{a_{f},b_{f}}\right\}$ and $A_{f}=D_{e}\cup\left\{{i}\right\}$. The proof of (E.3) is by recursion on $j$.
If $j=2$, then $f\in E_{1}$ so that $D_{f}=\varnothing$ while $D_{e}$ is a singleton, say $D_{e}=\left\{{d_{e}}\right\}$, and thus $f=\left\{{i,d_{e}}\right\}$. But then (E.3) follows from Assumption E.1(i) and Lemma E.3 below.
Let $j\geqslant 3$ and assume (E.3) has been shown for edges in $E_{1}\cup\cdots\cup E_{j-1}$. Let $k$ be the other node in $\mathcal{C}_{f}$ but $i$, that is, $\left\{{a_{f},b_{f}}\right\}=\mathcal{C}_{f}=\left\{{i,k}\right\}$. Then by (E.2) applied to $f$, we have
and thus
By the induction hypothesis, the two arguments of $C_{i|k;D_{f}}$ converge to $R_{i|D_{f}}(x_{i}|\bm{x}_{D_{f}})$ and $R_{k|D_{f}}(x_{k}|\bm{x}_{D_{f}})$, respectively. As $f\in E_{j-1}$ and $j\geqslant 3$, the pair copula density $c_{i,k;D_{f}}$ is continuous by Assumption E.1(ii). In view of Corollary E.5 below, $C_{i|k;D_{f}}$ is continuous in its two arguments too, and thus
By the recursive property (5.6), the last expression is equal to $R_{i|D_{f}\cup k}\left({x_{i}|\bm{x}_{D_{f}\cup k}}\right)$. As $D_{f}\cup k=D_{e}$, the convergence (E.3) follows. ∎
Lemma E.3.
The copula $C$ of $c$ and the tail copula $R$ of $r$ then satisfy
Proof of Lemma E.3.
But then
Lemma E.4.
Proof of Lemma E.4.
$\lim_{y^{\prime}\to y}\int_{I}\left|{f_{y^{\prime}}(x)-f_{y}(x)}\right|\,\mathrm{d}x=0.$ (E.4)
The first term on the right-hand side converges to zero by (E.4) and the second term vanishes asymptotically by the dominated convergence theorem. ∎
Corollary E.5.
Proof of Corollary E.5.
Apply Lemma E.4 to $f=c$ and $I=J=(0,1)$. ∎
Appendix F Additional numerical results
@@
F.1 Parametric relationships between tail dependence coefficients and bivariate tail copula parameters
@@
F.2 Additional simulation results to those in Section 8
F.2.1 Box-plots of dependence measure estimates with varying $(n,k)$

### v2

@@ B.5 Proofs for Section 6
Next, let $d\geqslant 4$ and assume the statement is true for regular vine sequence on $d-1$ elements. Let $e=(a_{e},b_{e};D_{e})=\left\{{f,g}\right\}$ be the single edge in $E_{d-1}$, with $f,g\in E_{d-2}$ labeled in such a way that $a_{e}\in\mathcal{C}_{f}$ and $b_{e}\in\mathcal{C}_{g}$. The regular vine sequence $\mathcal{V}$ induces regular vine sequences $\mathcal{V}_{f}$ and $\mathcal{V}_{g}$ on the node sets $A_{f}=[d]\setminus\left\{{b_{e}}\right\}$ and $A_{g}=[d]\setminus\left\{{a_{e}}\right\}$, both of which have $d-1$ elements; the edges $\bar{e}$ of the two regular vine sequences $\mathcal{V}_{f}$ and $\mathcal{V}_{g}$ are those of $\mathcal{V}$ that satisfy $A_{\bar{e}}\subseteq A_{f}$ or $A_{\bar{e}}\subseteq A_{g}$, respectively.
Let $j\in[d]$. If $j\neq b_{e}$, then we can apply the induction hypothesis to $\mathcal{V}_{f}$ to find a bijection $\sigma_{f,j}:\left\{{1,\ldots,d-1}\right\}\to A_{f}=[d]\setminus\left\{{b_{e}}\right\}$ such that (i) $\sigma_{f,j}(1)=j$ and (ii) $\sigma_{f,j}(k)$ belongs to the conditioned set of some edge $e_{k-1}$ in the $(k-1)$th tree of $\mathcal{V}_{f}$ as well as $\left\{{\sigma_{f,j}(i):i=1,\ldots,k}\right\}=A_{e_{k-1}}$, for all $k\in\left\{{2,\ldots,d-1}\right\}$. Extend $\sigma_{f,j}$ to a map $\sigma_{j}$ on $[d]$ by $\sigma_{j}(d)=b_{e}$. Then $\sigma_{j}$ meets all the requirements.
If $j=b_{e}$, then $j\neq a_{e}$, and we can apply the same argument to $\mathcal{V}_{g}$. ∎
@@ B.6 Proofs for Section 7
(i) By Definition 6.1, the density of $\bm{Z}$ is $\bm{z}\mapsto r(\bm{z})\operatorname{\mathds{1}}_{\mathbb{L}}(\bm{z})/R(\mathbb{L})$ with $\mathbb{L}=\left\{{\bm{x}:\min(\bm{z})<1}\right\}$. Let $K\subset[d]$ be non-empty and let $\bm{z}_{K}\in(0,\infty)^{K}$ be such that $\min\bm{z}_{K}<1$. The marginal density of $\bm{Z}_{K}=(Z_{k})_{k\in K}$ evaluated in $\bm{z}_{K}$ is
@@
Since $\min\bm{z}_{K}<1$, we also have $\min\bm{z}<1$ for $\bm{z}_{[d]\setminus K}\in(0,\infty)^{[d]\setminus K}$. As a consequence, the indicator function in the integral is always equal to one and, by Eq. (2.7), the density of $\bm{Z}_{K}$ evaluated in $\bm{z}_{K}$ simplifies to $r_{K}(\bm{z}_{K})/r(\mathbb{L})$.
Applying this identity to $K=J$ and $K=I\cup J$, we find, for $\bm{z}_{I}\in(0,\infty)^{I}$ and for $\bm{z}_{J}\in(0,\infty)^{J}$ such that $\min\bm{z}_{J}<1$ and $r_{J}(\bm{z}_{J})>0$, the conditional density of $\bm{Z}_{I}$ given $\bm{Z}_{J}=\bm{z}_{J}$ evaluated in $\bm{z}_{I}$:
@@
Appendix C Exponent measure densities as X-vines
@@
$c_{I;J}(\bm{u}_{I};1/\bm{y}_{J})=\frac{\lambda(\bm{y}_{I}|\bm{y}_{J})}{\prod_{i\in I}\lambda_{i|J}(y_{i}|\bm{y}_{J})}\quad\text{with}\quad y_{i}=\overline{\Lambda}_{i|J}^{-1}(u_{i}|\bm{y}_{J}).$ (C.1)
@@
Appendix D Estimating the copula parameters $\theta_{\mathcal{C}}$ in $T_{2},\ldots,T_{d-1}$.
Start with an edge $e=(a_{e},b_{e};D_{e})\in E_{2}$, i.e., such that $D_{e}$ is a singleton, $D_{e}=\left\{{d_{e}}\right\}$, say. Then $f=\left\{{a_{e},d_{e}}\right\}$ and $g=\left\{{b_{e},d_{e}}\right\}$ are edges in $E_{1}$. By Proposition 7.1(ii), $c_{a_{e},b_{e};D_{e}}(\,\cdot\,,\,\cdot\,;\theta_{e})$ is the density of the random pair $\left({R_{a_{e}|d_{e}}(Z_{a_{e}}|Z_{d_{e}};\theta_{f}),R_{b_{e}|d_{e}}(Z_{b_{e}}|Z_{d_{e}};\theta_{g})}\right)$. The tail copula density parameters $\theta_{f}$ and $\theta_{g}$ have been estimated in the previous step, yielding estimates $\hat{\theta}_{f}$ and $\hat{\theta}_{g}$. Define pseudo-observations $(\widehat{U}_{i,a_{e};D_{e}},\widehat{U}_{i,b_{e};D_{e}})$ for $i\in K_{D_{e}}=K_{d_{e}}$ from $c_{a_{e},b_{e};D_{e}}$ by $\widehat{U}_{i,a_{e};D_{e}}=R_{a_{e}|D_{e}}(\widehat{Z}_{i,a_{e}}\mid\widehat{Z}_{i,d_{e}};\;\hat{\theta}_{f})$ and $\widehat{U}_{i,b_{e};D_{e}}=R_{b_{e}|D_{e}}(\widehat{Z}_{i,b_{e}}\mid\widehat{Z}_{i,d_{e}};\;\hat{\theta}_{g})$. We estimate the parameter (vector) $\theta_{e}\in\Theta_{e}$ by maximising the pseudo-likelihood
$\mathcal{L}_{\mathcal{C},e}\left({\theta_{e};\;\left({\widehat{U}_{i,a_{e};D_{e}},\widehat{U}_{i,b_{e};D_{e}}}\right),i\in K_{D_{e}}}\right)=\prod_{i\in K_{D_{e}}}c_{a_{e},b_{e};D_{e}}\left({\widehat{U}_{i,a_{e};D_{e}},\widehat{U}_{i,b_{e};D_{e}};\,\theta_{e}}\right).$ (D.1)
@@
Other estimators are possible based on parametric relationships between dependence coefficients such as Kendall’s tau or Spearman’s rho and the parameters of corresponding bivariate copula densities (Czado, 2019).
Parameters (or parameter vectors) associated with edges $e=(a_{e},b_{e};D_{e})$ in $E_{j}$ for layers $j\in\left\{{3,\ldots,d-1}\right\}$ can be estimated similarly, proceeding recursively. Suppose that all parameters associated with edges in $E_{1:(j-1)}=E_{1}\cup\cdots\cup E_{j-1}$ have already been estimated; let $\theta(E_{1:(j-1)})$ denote the parameters and $\hat{\theta}(E_{1:(j-1)})$ their estimate. By Proposition 7.1(i), $c_{a_{e},b_{e};D_{e}}$ is the density of the random pair $\left({R_{a_{e}|D_{e}}\left({Z_{a_{e}}\mid\bm{Z}_{D_{e}};\;\theta(E_{1:j-1})}\right),\;R\left({Z_{b_{e}}\mid\bm{Z}_{D_{e}};\;\theta(E_{1:j-1})}\right)}\right)$. Define pseudo-observations $(\widehat{U}_{i,a_{e};D_{e}},\widehat{U}_{i,b_{e};D_{e}})$ for $i\in K_{D_{e}}$ from $c_{a_{e},b_{e};D_{e}}$ by
$\widehat{U}_{i,a_{e};D_{e}}=R_{a_{e}|D_{e}}\left({\widehat{Z}_{i,a_{e}}\mid\widehat{\bm{Z}}_{i,D_{e}};\;\hat{\theta}(E_{1:j-1})}\right),\quad\widehat{U}_{i,b_{e};D_{e}}=R_{b_{e}|D_{e}}\left({\widehat{Z}_{i,b_{e}}\mid\widehat{\bm{Z}}_{i,D_{e}};\,\hat{\theta}(E_{1:j-1})}\right).$
The conditional distribution functions $R_{a_{e}|D_{e}}$ and $R_{b_{e}|D_{e}}$ are calculated recursively via Eq. (5.7) with $e$ in that formula replaced by the edges $f,g\in E_{j-1}$ that are joined by $e$, that is, such that $e=\left\{{f,g}\right\}$. The parameter (vector) $\theta_{e}$ is then obtained by maximising the pseudo-likelihood (D.1).
Appendix E Tail dependence coefficients
E.1 Tail dependence coefficients for parametric families of tail copulas
@@
E.2 Tail dependence coefficients as goodness-of-fit measures
In Sections 8 and 9, we frequently use (pairwise) tail dependence coefficients to assess the goodness-of-fit of the estimated X-vine models. The aim of this section is to explain the difference between what we call model-based, empirical, and fitted tail dependence coefficients, and in particular, to explain how these are computed. First, recall the definition of the dependence coefficient $\chi_{J}$ for sets $J\subseteq[d]$ with two or more elements in Section 2. Note that
$\chi_{J}$ $=R\left({\left\{\bm{x}\in[0,\infty)^{d}:\max_{j\in J}x_{j}<1\right\}}\right)=R\left({\left\{\bm{x}\in[0,\infty)^{d}:\max_{j\in J}x_{j}<1\right\}\cap\mathbb{L}}\right)$
$=R\left({\mathbb{L}}\right)\operatorname{\mathsf{P}}\left({\max_{j\in J}Z_{j}<1}\right)=\operatorname{\mathsf{P}}\left({\max_{j\in J\setminus k}Z_{j}<1\mid Z_{k}<1}\right),$
since $\operatorname{\mathsf{P}}\left({Z_{k}<1}\right)=1/R(\mathbb{L})$ for any $k\in J$. Empirical estimates of $\chi_{J}$ will be based on the right-most expression in above equation.
E.2.1 Model-based tail dependence coefficients
Model-based $\chi$’s, used in Section 8, are those corresponding to the true X-vine specification. The parametric expressions given in Section E.1 can only be used for pairs of variables connected by an edge in the first tree of the X-vine, i.e., for $\chi_{ab}$ with $\{a,b\}\in E_{1}$. More generally, model-based $\chi_{J}$’s for sets $J\subseteq[d]$ with two or more elements will have to be calculated via Monte Carlo simulation:
  1. 1.
    Simulate a sample of size $N$ from the inverted multivariate Pareto distribution associated with the true X-vine specification, $\bm{Z}_{1},\ldots,\bm{Z}_{N}$. In Section 8, we used $N=5000$.
  2. 2.
    Use this sample to estimate $\chi_{J}$ for non-empty $J\subseteq[d]$ as the average of the $|J|$ ratios
    $\frac{|\bigcap_{j\in J}Q_{j}|}{|Q_{k}|},\qquad k\in J,$ (E.1)
    where $Q_{j}=\left\{{i=1,\ldots,N:Z_{i,j}<1}\right\}$ for $j\in[d]$.
E.2.2 Empirical tail dependence coefficients
In Section 9, we calculate the empirical tail dependence coefficients for the large flight delays data $\hat{\bm{x}}_{1},\ldots,\hat{\bm{x}}_{n}$ as follows.
  1. 1.
    Apply the rank transformation of Eq. (7.1) to obtain a sample $\hat{\bm{u}}_{1},\ldots,\hat{\bm{u}}_{n}$.
  2. 2.
    For non-empty $J\subseteq[d]$, estimate $\chi_{J}$ by
    $\widehat{\chi}_{J}=\frac{1}{k}\sum_{i=1}^{n}\prod_{j\in J}\operatorname{\mathds{1}}\left\{{\hat{u}_{i,j}<k/n}\right\}.$
Using instead (E.1) applied to the sample $\hat{\bm{z}}_{i}=(n/k)\hat{\bm{u}}_{i}$ for $i=1,\ldots,n$ leads to the same estimate as $\widehat{\chi}_{J}$. Indeed, replacing $Z_{i,j}$ by $\hat{z}_{i,j}$ in the definition of $Q_{j}$ gives the set $K_{j}=\{i=1,\ldots,n:\hat{u}_{i,j}<1\}$ and $|K_{j}|=k$ for all $j\in[d]$.
E.2.3 Fitted tail dependence coefficients
Sections 8 and 9 both use fitted tail dependence coefficients. In the simulation study in Section 8, these are calculated as the model-based tail dependence coefficients described above; the only difference is that in step 1, we simulate a sample of size $N=5000$ from the fitted X-vine model. However, in Section 9, the fitted tail dependence coefficients are calculated as follows:
  1. 1.
    Simulate a sample of size $N=3600$ (approximately the same sample size as that of the data) from the inverted multivariate Pareto distribution associated with the fitted X-vine model, $\bm{Z}_{1},\ldots,\bm{Z}_{N}$.
  2. 2.
    Apply the procedure outlined in the paragraph “Empirical tail dependence coefficients” to the multivariate Pareto sample $\bm{Z}_{1},\ldots,\bm{Z}_{N}$.
The adjusted sample size of $N=3600$ and our treatment of the multivariate Pareto sample as if stemming from an unknown distribution ensure that the fitted tail dependence coefficients are comparable to the empirical ones in terms of bias.
Appendix F Additional numerical results
F.1 Additional simulation results to those in Section 8
F.1.1 Box-plots of dependence measure estimates based on pre-specified bivariate parametric families for each edge
Similar to Fig. LABEL:fig:boxplot_Fam, but this time assuming the same vine structure and the pre-specified bivariate parametric families in the X-vine specification shown in Fig. LABEL:fig:5d_Xvine_a, we create box-plots of dependence measure estimates in Fig. F.2. As expected, we see lower variability and less bias, especially in deeper trees (e.g. $T_{3}$ and $T_{4}$).
Figure F.2: Box-plots ($\blacksquare$ for $T_{1}$blacksquare for $T_{2}$blacksquare for $T_{3}$blacksquare for $T_{4}$) of dependence measure estimates from specified bivariate parametric families for each edge across the four trees in the X-vine specification in Fig. LABEL:fig:5d_Xvine_a.
F.1.2 Box-plots of dependence measure estimates with varying $(n,k)$